12. Find the potential drop along a copper wire 500 m long and 2 mm in diameter if the current in it is equal to 2 A


1
Expert's answer
2022-01-30T14:18:41-0500

Solution

Length of wire l=500m

Radius R=1mm

So area "A=\pi r^2=3.14\times1mm\times1mm"

"A=3.14\times10^{-6}m^2"

Resistivity of copper wire

"\rho=1.68 * 10^{-8} Ω -m"

So potential difference of wire

"V=IR\\=\frac{I\rho l}{A}"

Putting all values

"V=\frac{2\times1.68 * 10^{-8}\times500 }{3.14\times10^{-6}}\\=5.35Volts"